OpenStudy (anonymous):

Five times the greater the greater of two consecutive integers is equal to 21 more than the lesser. Find the 2 Integers

jimthompson5910 (jim_thompson5910):

5(x+1)=x+21 5x+5=x+21 5x-x=21-5 4x=16 x=16/4 x=4 So the first number is 4 and the second number is 5